# Stacked Intelligent Metasurface-Enhanced Wideband Multiuser MIMO OFDM-IM Communications

### v1

@@ IV-A Motivation and Design Principle
    Double unrolling (layer index $l$ and iteration $t$) matches the $L$-layer cascade and the $T$ training iterations;
@@ IV-B UPGD Network Architecture
Figure 3: UPGD-Net for jointly optimizing the phase shifts of the $L$-layer SIM over $T$ projected-gradient iterations. Columns denote successive iterations t=0:T1t=0:T-1, and rows correspond to SIM layers $l=1:L$. In each cell, the analytic gradient $\nabla$ is scaled by the trainable step size $\eta^{(t)}$, added to the current phase-shift tensors, and then projected to $0,2\pi)$. A vertical coupling bus delivers the worst-link SINR gradient, computed from the cascade precoder of all layers, to every $\nabla$-block within the same iteration. All modules are model-driven and fully differentiable.
@@ IV-B1 Forward Propagation:
The inputs of the proposed UPGD-Net are: (i) the loss function matrix $\mathcal{H}=\{\mathbf{H},\mathbf{Z},\mathbf{W},\mathbf{p}\}$ for SINR calculation, using the channel matrix $\mathbf{H}$, subcarrier allocation $\mathbf{Z}$, SIM transmission matrix $\mathbf{W}$, and the power allocation matrix $\mathbf{p}$; (ii) the initialized phase-shift tensor ${\bm{\theta}}^{(0)}=\{{\bm{\theta}}^{1(0)},\ldots,{\bm{\theta}}^{L(0)}\}$, where ${\bm{\theta}}^{l(0)}={\{\theta_{m}^{l(0)}\}}_{m\in\mathcal{M}}$. In each iteration $t$, the following computations are performed:
@@ IV-B2 Backward Propagation:
    Model-driven structure: The architecture strictly follows the iterative PGD structure, leveraging domain knowledge from fully-analog beamforming and wave-based SIM optimization.
@@
    Data-driven learning: The step sizes and update trajectories are learned from data, adapting to practical channel variations and transmission scenarios.
@@ V-A Simulation Configurations
The system operates at a center frequency $f_{0}\!=\!28$ GHz over a bandwidth $B_{w}\!=\!60$ MHz, and the number of subcarriers is set as $N_{c}\!=\!16$. The BS is mounted at a height of $10$ m, while the $K$ single-antenna UEs are placed $250$ m away on a line-of-sight street, spaced $30$ m apart. Following 3GPP TR 38.901 guidelines for mmWave outdoor micro-cells [57], we set $P_{k}\!=\!10$ dominant taps with delays. The length of CP is set to $N_{cp}\!=\!8$ samples (200 ns) for ensuring that $N_{cp}$ is greater than the maximum delay $\tau_{p-\rm{max}}$, thus mitigating inter symbol interference. The thickness of SIM is set as $D_{m}\!=\!0.05$ m, and the spacing between the layers is set as $d_{m}\!=\!D_{m}/L$. The parameters of SIM are configured with S=K=4S\!=\!K\!=\!4, M=100M\!=\!100, L=7L\!=\!7, $r_{m}=c/(2f_{0})$, and $S_{m}\!=\!c^{2}/(2f_{0})^{2}$. The transmit power $P_{t}$ is set to 10 dBm and the antenna gains of the BS and UEs are set to $3$ dBi and 0 dBi, respectively. The noise power spectral density is -174 dBm/Hz. The BS power allocation matrix $\mathbf{p}$ is optimized based on the iterative water-filling algorithm [41]. For active subcarrier index selection, $L_{b}\!=\!4$, N=4N\!=\!4, and V=2V\!=\!2 are firstly configured in Table II, where $q_{1}\!=\!2$ bits is used to determine the indices of the two active subcarriers out of four subcarriers.
@@
UPGD-Net implements the double-unfold grid of Fig. 3: each column corresponds to one projected-gradient $\bm{\theta}$-update block. The step-size schedules $\{\eta^{(t)}\}_{t=0}^{T-1}$ constitute the only trainable parameters with T=30T\!=\!30 unfolded stages. The learning database comprises 5,000 pairs of independently loss function matrices $\mathcal{H}$ for different bitstream data, while an 80%:20% split is used for training and validation. Offline optimization relies on the Adam algorithm with a learning rate of $1\times 10^{-3}$, a batch size of 64, and 120 training epochs.
@@ V-B Convergence of the Proposed UPGD-Net
Fig. 4 confirms the superiority of the proposed UPGD-Net design. The solid orange curve is the mean performance of the proposed UPGD-Net with learned step sizes, averaged over 1,000 independent phase-shift realizations and randomly generated OFDM-IM activation matrices $\mathbf{Z}$. The surrounding shaded area shows the corresponding $\pm 1~\!\sigma$ range, i.e., the region bounded by the 16-th and 84-th percentiles of the 1,000 trajectories, illustrating the spread induced by bitstream variability. The dashed blue curve depicts conventional PGD with an a priori tuned, fixed step, while the dotted black line marks the CVX upper bound obtained by an offline algorithm in [41]. Starting from the same initial condition of 0, UPGD-Net drives the loss below -$4.5\times 10^{-8}$ within 12 unfolded stages and reaches -$5\times 10^{-8}$ after 25 stages, less than $0.2\times 10^{-8}$ away from the CVX global optimum. In contrast, the fixed-step PGD requires almost 50 iterations to attain a comparable level, demonstrating a convergence speed-up of 2.5 times. The red band further highlights the generalization capability: although every realization employs a different index bitstream, with a different subcarrier activation matrix $\mathbf{Z}$, the learned step-size schedule keeps all trajectories tightly clustered. This behavior underpins the data-driven component of the network, while the strictly PGD-based layer structure preserves the model-driven monotonic descent, ensuring that no sample ever diverges.
@@
Fig. 5 illustrates how the number of transmit metasurface layers $L$ influences the optimization carried out by the proposed UPGD-Net. With a fixed per-layer aperture with the number of meta-atoms M=100M\!=\!100, stacking more layers provides additional DoF, so the proposed UPGD-Net converges faster and reaches a lower loss from increasing L=3L\!=\!3 to L=7L\!=\!7. When $L$ increases further to 9, though the performance converges fast in the initial iteration, it slightly degrades in the final part. This is because neighboring layers become strongly electromagnetically coupled, so each additional layer under a fixed thickness manipulates the wavefront less effectively when exceeding a certain threshold. A larger $L$ also enlarges the two-dimensional unfolding grid, increasing per-iteration complexity, as every projected-gradient update has to account for the phase-shift tensors of all preceding layers. Under the present hardware budget, L=7L\!=\!7 therefore emerges as the empirical sweet spot, striking a trade-off between worst-link SINR optimization and implementation cost. Each curve represents the ensemble mean over 1,000 random OFDM-IM activation matrices, while the surrounding shaded area marks the corresponding $\pm 1~\sigma$ interval. These results confirm that the double-unrolled schedule on training iteration and metasurface layer generalizes well across various bit-stream variability, underscoring the robustness of the proposed model-/data-hybrid design.
@@ V-C End-to-end System Performance Evaluation
Figure 6: Average BER versus total transmit power for the proposed SIM-enhanced OFDM-IM and three baseline schemes.
@@
Figure 7: Sum rate versus total transmit power for the proposed SIM-enhanced OFDM-IM and two baseline schemes.
Fig. 7 plots the energy-normalized sum rate versus the total transmit power for the proposed SIM-enhanced OFDM-IM, an OFDM-IM with digital ZF baseline, and classical OFDM with digital ZF. Across the entire transmit power sweep, the proposed SIM-enhanced OFDM-IM curve lies 4040\, $80\%$ above its digital ZF counterparts. The gain is attributed to the large-aperture SIM, which adds a wave-domain boost to every active subcarrier. In addition, digital ZF requires energy normalization to satisfy the total power constraint, while wave-domain precoding via SIM does not amplify coefficients and redistributes energy more effectively across space. Moreover, classical OFDM distributes the same transmit power over all sixteen subcarriers; the per-tone SNR is therefore halved w.r.t. OFDM-IM, resulting in the lowest sum-rate. In contrast, index modulation concentrates power on the $K=2$ active tones while injecting additional index bits through the allocation matrix $\mathbf{Z}$. Together with the previous BER results in Fig. 6, this joint exploitation of SIM aperture gain and index modulation diversity yields the highest sum rate for a given power budget, confirming that the proposed transceiver architecture offers superior power efficiency at only a modest increase in detection complexity.
@@ V-D Trade-off of the Index Selection in OFDM-IM
V-E Comparison with Existing Transmission Technologies
((a)) TX1 antenna
((b)) TX2 antenna
((c)) TX3 antenna
((d)) TX4 antenna
Figure 9: Instantaneous envelope magnitudes of time-domain waveforms at four TX antennas with RF chains for three schemes, where all cases employ BPSK modulation and the same average transmit power.
Finally, Fig. 9 illustrates the PAPR of the time-domain waveforms at four TX antennas with RF chains for three transmission schemes: (i) SIM-enhanced OFDM-IM with wave-domain ZF, (ii) OFDM-IM with conventional digital ZF, and (iii) classical full-subcarrier BPSK-OFDM with digital ZF. For clarity, the maximum and average amplitudes are indicated by dashed lines in each subfigure, such that the vertical gap directly visualizes the PAPR. It is evident that the proposed SIM-enhanced OFDM-IM consistently achieves the lowest PAPR across all TX antennas, with values around 3.53.5\, 3.9\,3.9 dB. In contrast, conventional OFDM-IM yields moderately reduced PAPR levels of 55\, 6\,6 dB, while traditional OFDM exhibits the largest fluctuations with PAPR close to $8.3$ dB.
These results confirm that by jointly exploiting index modulation and wave-domain processing enabled by SIM, the proposed transceiver significantly suppresses envelope peaks, thereby mitigating the intrinsic high-PAPR drawback of OFDM systems. Specifically, the digital baseband directly feeds raw signals into the RF chains, which are subsequently processed by the SIM. This fully-analog wave-domain ZF precoding avoids symbol-level superposition across data streams, while the superposition caused by digital ZF precoding across four data streams of OFDM-IM and OFDM systems further increases constructive peak values. Additionally, index modulation activates only a subset of subcarriers within each subblock, which reduces the instantaneous signal power variance in the time domain, which mitigates large signal spikes and thus suppresses the PAPR. By contrast, the conventional full-tone OFDM with digital ZF generates the most fluctuating waveforms, since every subcarrier is active and carries independent BPSK symbols. Moreover, the superposition caused by digital ZF precoding across four data streams of OFDM-IM and OFDM systems further increases peak values. Overall, the reduction of 22\, 5\,5 dB by the proposed SIM-enhanced architecture over the baselines translates into lower back-off requirements for the power amplifier, improved energy efficiency, and enhanced feasibility of practical deployment.

### v2

@@ IV-A Motivation and Design Principle
    Coupled physical-algorithmic unrolling: The nested computational graph aligns the algorithmic iterations $t$ with the physical wave propagation through layer indices $l$, enforcing physics-compliant gradient back-propagation;
@@ IV-B UPGD Network Architecture
Figure 3: UPGD-Net for optimizing the phase shifts of the $L$-layer SIM over $T$ projected-gradient iterations. Columns denote successive iterations t=0:T1t=0:T-1, and rows correspond to SIM layers $l=1:L$. In each cell, the analytic gradient $\nabla$ is scaled by the trainable step size $\eta^{(t)}$, added to the current phase-shift tensors, and then projected to $0,2\pi)$. A vertical coupling bus delivers the worst-link SINR gradient, computed from the cascade precoder of all layers, to every $\nabla$-block within the same iteration. All modules are model-driven and fully differentiable.
@@ IV-B1 Forward Propagation:
The inputs of the proposed UPGD-Net are: (i) the loss function matrix $\mathcal{H}=\{\mathbf{H},\mathbf{Z},\mathbf{W},\mathbf{p}\}$ for SINR calculation, using the channel matrix $\mathbf{H}$, activation matrix $\mathbf{Z}$, SIM transmission matrix $\mathbf{W}=\{W_{n}^{(l)}\}$, and the power allocation matrix $\mathbf{p}$; (ii) the initialized phase-shift tensor ${\bm{\Theta}}^{(0)}=\{{\bm{\theta}}^{1(0)},\ldots,{\bm{\theta}}^{L(0)}\}$, where ${\bm{\theta}}^{l(0)}={\{\theta_{m}^{l(0)}\}}_{m\in\mathcal{M}}$. In each iteration $t$, the following computations are performed:
@@ IV-B2 Backward Propagation:
The end-to-end training with per-iteration shared steps is demonstrated in Algorithm 1. Each forward pass unrolls $T$ projected-gradient steps, and the gradient is recomputed in the next iteration using the previous one. After $T$ steps, the final test loss $\min_{k,\ell,n}\,\gamma_{k,\ell,n}^{(T)}$ is backpropagated through the unrolled structure to update step sizes, while phase shifts are re-optimized per frame by the unrolled solver and are not learned across data. At inference time the epoch loop and backprop are disabled, we freeze $\{\eta^{(t)}\}_{t=0}^{T-1}$ and run the forward unroll once, which yields a near-optimal phase shift tensor in $T$ lightweight updates.
Initialization: Phase shifts $\bm{\Theta}^{(0)}$, channels $\mathbf{H}$, per-layer transmissions $\mathbf{W}$, activation matrix $\mathbf{Z}$, power allocation matrix $\mathbf{p}$, layers $L$, training epoch $E$, iterations $T$, trainable step sizes $\{\eta^{(t)}\}_{t=0}^{T-1}$.
1:for $epoch=1$ to $E$ do
2:  for $t=0$ to T1T-1 do
3:   // Form cascade with current phases $\bm{\Theta}^{(t)}$
4:   For each subcarrier $n$, form $\mathbf{G}_{n}^{(t)}=\prod_{l=L}^{1}\mathbf{\Phi}^{l(t)}\mathbf{W}_{n}^{l}$, with $\mathbf{\Phi}^{l(t)}=\mathrm{diag}\!\big(e^{j\theta^{l(t)}_{1}},\ldots,e^{j\theta^{l(t)}_{M}}\big)$.
5:   Compute instantaneous SINR $\gamma_{k,\ell,n}^{(t)}$ on all active triplets ($Z_{k,\ell,n}{=}1$) via (12).
6:   Find worst-link $(k^{*},\!\ell^{*},\!n^{*})$ and set $\mathcal{L}^{(t)}$ via (28).
7:   Compute $\mathcal{G}^{(t)}=\left\{\frac{\partial\gamma_{k^{*},\ell^{*},n^{*}}^{(t)}}{\partial\theta_{m}^{l(t)}}\right\}_{l,m}$ via (29), and broadcast $\mathcal{G}^{(t)}$ to all $L$ matesurface layers.
8:   // Layer-wise projected update with shared step $\eta^{(t)}$
9:   for $l=1$ to $L$ do
10:    for $m=1$ to $M$ do
11:     Update phase shift $\theta_{m}^{l(t+1)}\leftarrow{\text{ Proj}}_{[0,2\pi)}\big(\theta_{m}^{l(t)}-\,\eta^{(t)}\cdot\mathcal{G}^{(t)}(l,m)\big)$.
12:    end for
13:   end for
14:  end for
15:  Final loss for training: $\mathcal{L}_{\text{total}}\leftarrow-\min_{k,\ell,n}\gamma_{k,\ell,n}^{(T)}$.
16:  for $t=0$ to T1T-1 do
17:   // Backpropagate $\mathcal{L}_{\text{total}}$ through the unfolded PGD
18:   Compute $\frac{\partial\mathcal{L}_{\text{total}}}{\partial\eta^{(t)}}$ by chain rule (32).
19:   Update step size $\eta^{(t)}\leftarrow\text{Adam}\big(\eta^{(t)},\frac{\partial\mathcal{L}_{\text{total}}}{\partial\eta^{(t)}}\big)$.
20:  end for
21:end for
22:Output: Learned step-size schedule $\{\eta^{(t)}\}$ and final phases $\bm{\Theta}^{(T)}$.
Algorithm 1 UPGD-Net: End-to-end training with per-iteration shared steps
@@
    Model-driven structure: The architecture strictly follows the iterative PGD structure with monotone descent, leveraging domain knowledge from fully-analog beamforming and wave-based SIM optimization.
@@
    Data-driven learning: The step sizes and update schedule are learned from data, adapting to practical channel variations and transmission scenarios.
@@ V-A Simulation Configurations
The system operates at a center frequency $f_{0}\!=\!28$ GHz over a bandwidth $B_{w}\!=\!60$ MHz, and the number of subcarriers is set as $N_{c}\!=\!16$. The BS is mounted at a height of $10$ m, while the $K$ single-antenna UEs are placed $250$ m away on a line-of-sight street, spaced $30$ m apart. Following 3GPP TR 38.901 guidelines for mmWave outdoor micro-cells [59], we set $P_{k}\!=\!10$ dominant taps with delays. The length of CP is set to $N_{cp}\!=\!8$ samples (200 ns) for ensuring that $N_{cp}$ is greater than the maximum delay $\tau_{p-\rm{max}}$, thus mitigating inter symbol interference. The thickness of SIM is set as $D_{m}\!=\!0.05$ m, and the spacing between the layers is set as $d_{m}\!=\!D_{m}/L$. The parameters of SIM are configured with S=K=4S\!=\!K\!=\!4, M=100M\!=\!100, L=7L\!=\!7, $r_{m}=c/(2f_{0})$, and $S_{m}\!=\!c^{2}/(2f_{0})^{2}$. The transmit power $P_{t}$ is set to 10 dBm and the antenna gains of the BS and UEs are set to $5$ dBi and 0 dBi, respectively. The noise power spectral density is -174 dBm/Hz. The BS power allocation matrix $\mathbf{p}$ is optimized based on the iterative water-filling algorithm [41]. For active subcarrier index selection, $L_{b}\!=\!4$, N=4N\!=\!4, and V=2V\!=\!2 are firstly configured in Table II, where $q_{1}\!=\!2$ bits is used to determine the indices of the two active subcarriers out of four subcarriers.
@@
UPGD-Net implements the physical-algorithmic unrolling in Fig. 3: each column corresponds to one projected-gradient $\bm{\theta}$-update block. The step-size schedules $\{\eta^{(t)}\}_{t=0}^{T-1}$ constitute the only trainable parameters with T=30T\!=\!30 unfolded stages. The learning database comprises 5,000 pairs of independently loss function matrices $\mathcal{H}$ for different bitstream data, while an 80%:20% split is used for training and validation. Offline optimization relies on the Adam algorithm with a learning rate of $1\times 10^{-3}$, a batch size of 64, and 120 training epochs.
@@ V-B Convergence of the Proposed UPGD-Net
Fig. 4 confirms the superiority of the proposed UPGD-Net design. The solid orange curve is the mean performance of the proposed UPGD-Net with learned step sizes, averaged over 1,000 independent phase-shift realizations and randomly generated OFDM-IM activation matrices $\mathbf{Z}$. The surrounding shaded area shows the corresponding $\pm 1~\!\sigma$ range, i.e., the region bounded by the 16-th and 84-th percentiles of the 1,000 trajectories, illustrating the spread induced by bitstream variability. The dashed blue curve depicts conventional PGD with an a priori tuned, fixed step, while the dotted black line marks the CVX upper bound obtained by an offline block coordinate descent (BCD) algorithm in [41]. Starting from the same initial condition of 0, UPGD-Net drives the loss below -$4.5\times 10^{-2}$ within 12 unfolded stages and reaches -$5\times 10^{-2}$ after 25 stages, less than $0.2\times 10^{-2}$ away from the CVX global optimum. In contrast, the fixed-step PGD requires almost 50 iterations to attain a comparable level, demonstrating a convergence speed-up of 2.5 times. The red band further highlights the generalization capability: although every realization employs a different index bitstream, with a different subcarrier activation matrix $\mathbf{Z}$, the learned step-size schedule keeps all trajectories tightly clustered. This behavior underpins the data-driven component of the network, while the strictly PGD-based layer structure preserves the model-driven monotonic descent, ensuring that no sample ever diverges.
@@
Fig. 5 illustrates how the number of transmit metasurface layers $L$ influences the optimization carried out by the proposed UPGD-Net. With a fixed per-layer aperture with the number of meta-atoms M=100M\!=\!100, stacking more layers provides additional DoF, so the proposed UPGD-Net converges faster and reaches a lower loss from increasing L=3L\!=\!3 to L=7L\!=\!7. When $L$ increases further to 9, though the performance converges fast in the initial iteration, it slightly degrades in the final part. This is because neighboring layers become strongly electromagnetically coupled, so each additional layer under a fixed thickness manipulates the wavefront less effectively when exceeding a certain threshold. A larger $L$ also enlarges the two-dimensional unfolding grid, increasing per-iteration complexity, as every projected-gradient update has to account for the phase-shift tensors of all preceding layers. Under the present hardware budget, L=7L\!=\!7 therefore emerges as the empirical sweet spot, striking a trade-off between worst-link SINR optimization and implementation cost. Each curve represents the ensemble mean over 1,000 random OFDM-IM activation matrices, while the surrounding shaded area marks the corresponding $\pm 1~\sigma$ interval. These results confirm that the coupled physical-algorithmic unrolling schedule on training iteration and metasurface layer generalizes well across various bit-stream variability, underscoring the robustness of the proposed model-/data-hybrid design.
@@ V-C End-to-end System Performance Evaluation
Figure 6: Average BER versus total transmit power for the proposed SIM-enhanced OFDM-IM and two baselines: digital-ZF OFDM-IM and digital-ZF full-tone OFDM.
@@
Figure 7: Sum rate versus total transmit power for the proposed SIM-enhanced OFDM-IM and three baselines: digital-ZF OFDM-IM, digital-ZF full-tone OFDM, and hybrid-precoding OFDM.
Fig. 7 depicts the energy-normalized sum rate versus the total transmit power for four schemes: the proposed SIM-enhanced OFDM-IM, a digital-ZF OFDM-IM baseline, a classical full-tone OFDM with digital ZF, and a large-aperture hybrid precoding OFDM baseline. To ensure a fair comparison in terms of active hardware complexity, all schemes are constrained to the same RF budget of $N_{RF}\!=\!4$ (serving K=4K\!=\!4 users). Across the entire transmit power sweep, the proposed SIM-enhanced OFDM-IM curve lies 4040\, $80\%$ above its digital ZF counterparts. The gain is attributed to the large-aperture SIM (M=100M\!=\!100 controllable metasurface phase shifts per layer), which adds a wave-domain boost to every active subcarrier. Unlike digital ZF, which often suffers from a severe power normalization penalty to satisfy the total power constraint, the SIM redistributes signal energy more effectively via diffractive propagation. It is also worth noting that the two digital-ZF baselines exhibit closely matched sum-rate curves. This is because, under the BPSK modulation scheme, the spectral efficiency loss of IM is compensated by the power concentration on active subcarriers, resulting in a comparable overall payload to that of classical full-tone OFDM under the same system budget.
In Fig. 7, we further include a hybrid precoding OFDM benchmark that uses a conventional fully connected phase-shifter network to drive a 100-element antenna array with the same 4 RF chains. As expected, the larger antenna aperture improves the achievable sum rate over the $N_{TX}=4$ digital baselines, especially in the low-SNR regime. Nevertheless, the proposed SIM-enhanced OFDM-IM still offers a substantial gain, highlighting that the multi-layer structure can exploit a large passive aperture more effectively than a conventional hybrid precoder under the same RF-chain constraint. In practice, the hybrid precoder relies on an active aperture requiring 100 power amplifiers and hundreds of active phase shifters to drive the antenna array, while the SIM achieves equivalent performance using a quasi-passive multilayer structure with only 4 active RF chains. Consequently, the proposed SIM architecture offers a decisive advantage in terms of power consumption and hardware cost compared to conventional active phased arrays. This benefit is even more significant when compared against a fully-digital massive-MIMO benchmark, which would necessitate scaling the number of RF chains linearly with the number of antennas (i.e., $N_{RF}\!=\!N_{TX}\!=\!100$).
To address practical concerns regarding imperfect CSI at the BS, we evaluate the robustness of the proposed BER-driven UPGD-Net design. Fig. 8 depicts the average BER performance versus the normalized CSI error variance, $\xi\in[0,0.1]$, for both the proposed SIM-enhanced OFDM-IM scheme and the digital ZF baseline. As expected, the BER for both schemes degrades with increasing CSI error due to the mismatch between the true and estimated channels. However, the degradation trends differ significantly. The digital ZF baseline exhibits a rapid performance deterioration as the error variance grows. This behavior stems from the inherent sensitivity of ZF precoding to channel estimation errors, where matrix inversion amplifies estimation noise, resulting in severe residual inter-user interference. In contrast, the proposed SIM-based approach demonstrates superior robustness with a much more graceful BER degradation, thereby widening the performance gap over the ZF baseline at moderate-to-high error levels. This resilience is attributed to the gradient-based wave-domain optimization inherent in the SIM architecture and the unfolded UPGD-Net. The UPGD-Net leverages gradient-based wave-domain optimization to enhance the effective SINR margin in a BER-driven manner. Consequently, the learned beam pattern maintains its effectiveness even under CSI perturbations, validating the practical applicability of the proposed design in realistic TDD deployments.
Figure 8: Average BER versus normalized CSI error variance for the proposed SIM-enhanced OFDM-IM structure with UPGD-Net and the digital ZF baseline.
@@ V-D Trade-off of the Index Selection in OFDM-IM
V-E PAPR Reduction and RF Efficiency Analysis
((a))
((b))
Figure 10: PAPR characteristics and time-domain envelope analysis: (a) Instantaneous envelope magnitude $|x|$ at the TX antenna. (b) Statistical CCDF of the PAPR over randomly generated OFDM symbols and all TX antennas.
Finally, Fig. 10 characterizes the PAPR of the three considered schemes: (i) SIM–enhanced OFDM–IM with wave–domain ZF, (ii) OFDM–IM with conventional digital ZF, and (iii) classical full–subcarrier OFDM with digital ZF. Fig. 10(a) first illustrates the instantaneous envelope $|x(t)|$ at a representative TX antenna for one OFDM symbol. The dynamic range between the maximum (red dashed) and average (cyan dashed) amplitudes directly reflects the PAPR level. The SIM–enhanced OFDM–IM waveform exhibits the flattest envelope, whereas OFDM–IM with digital ZF shows moderate peaks, and conventional OFDM with digital ZF exhibits the most severe fluctuations.
To provide a statistical assessment, Fig. 10(b) plots the complementary cumulative distribution function (CCDF) of the PAPR, evaluated over $10^{4}$ randomly generated OFDM symbols and all four TX antennas, with all schemes normalized to the same average transmit power. The proposed SIM-enhanced OFDM-IM clearly achieves the most favorable PAPR statistics: its CCDF curve lies consistently to the left of the two digital baselines, and at a CCDF level of $10^{-3}$, the required back-off is approximately $6$ dB, compared with about 8899 dB for OFDM–IM with digital ZF and roughly 9.59.51010 dB for classical OFDM with digital ZF. This gain originates from two mechanisms: First, beamforming and multiuser interference suppression are realized in the wave domain by the SIM, so each power amplifier amplifies a single-stream OFDM-IM waveform rather than a digitally superimposed multi-stream signal, thereby avoiding the mixing-induced PAPR penalty of conventional MIMO precoding. Second, index modulation activates only a subset of subcarriers in each subblock, which reduces the instantaneous power variance relative to full-tone OFDM. Overall, the observed 2255 dB PAPR reduction of the SIM-enhanced architecture over digital ZF baselines directly translates into more relaxed power-amplifier back-off requirements and improved energy efficiency in practical implementations.